OpenStudy (anonymous):

Select the compound in which chlorine is assigned the oxidation number +5. a) HClO4 b) HC1O2 c) HCl d) NaClO3 e) HClO Please help, I don't know what to do!

OpenStudy (ivancsc1996):

Well you must add up all the individual oxidation numbers using chlorine as five, and the one that adds up to zero is the answer since it is the stable one that can exist. a)\[HClO _{4} \rightarrow H:1 + Cl:5 + 4(O:-2) =-2 \]b)\[HClO _{2} \rightarrow H:1 + Cl:5 + 2(O:-2)=2\]c) \[HCl \rightarrow H:1 + Cl:5=6\]d) \[NaClO _{3} \rightarrow Na:1+Cl:5+3(O:-2)=0\]e) \[HClO \rightarrow H:1 + Cl: 5 + O:-2 = 4\]The answer is d). Now I don't want to confuse you, some of the other ones can exist but they don't have chlorine with the oxidation number five, they insted use other oxidation numbers such as: \[HCl \rightarrow H ^{1}Cl ^{-1} \rightarrow H:1 + Cl: -1=0\]